OpenStudy (anonymous):

solve for x and y? 2y-7x=-9, 4x+6=y

OpenStudy (anonymous):

4x+6=y put y in other equation 2*(4x+6)-7x=-9 8x+12-7x=-9 now solve it

OpenStudy (anonymous):

thank you!!!